divide 150 into three parts such that the second number is five-sixths the first and the third number is four-fifths the second.
Answer by amoresroy(361)

Let x = the first number
5x/6= the second number
(4/5)(5x/6) = the third number= 2x/3
x+5x/6+2x/3 = 150
Multiply both sides by 6 to get
6x+5x+4x = 900
x = 900/15
x = 60 (first number)
2nd number:
5*60/6 = 50
3rd number:
4*50/5 = 40
The numbers are 60, 50, and 40.